7 + 3(-2) - 4/2 +1 =
Answer by benbjlin(6)

You must solve multiplication and division problems first before addition and subtraction. Always solve left to right.
1. 7 + 3*(-2) - 4/2 + 1 = ?; 3*(-2)=(-6), 4/2=2
2. 7 + (-6) - 2 + 1 = ?; 7+(-6)=7-6=1
3. 1 - 2 + 1 = ?; 1-2=(-1), (-1)+1=0
4. 7 + 3*(-2) - 4/2 + 1 = 0
5. Ans=0